Jiffy Bros., Inc. has total capital (V=D+E) of $10.8 million. Jiffy also has 4,139 bonds outstanding that currently sell at 103% of face value ($1,000). If Jiffy has 34,718 shares of stock what is We, the weight of equity in Jiffy's capital structure? (please show me the process!Thank you!)
Answer #1

Calculation of weight of Equity

Total book value of debt = 4139*1030=42,63,170

So Total Equity = 108,00,000-42,63,170=65,36,830

So weight of equity = 65,36,830/108,00,000=60.53%